state which is greater 
 
(8-9)×10 or 8-9×10       ???????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????

Grade:8

1 Answers

Shahid
7 years ago
Dear  Samiksha , 
Here , You must use the “BODMAS” Rule.
BODMAS stands for  brackets, order, division, multiplication, addition and subtraction.
 
Therefore
(8-9) x 10 = (-1) x 10  = (-10)
8 – 9 x 10   =  8 – 90  = (-82) 
 
Therefore, (8-9) x 10 is Greater.
 
Hope It Helps.
With Regards.
Shahid.